OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

I got the same =)

OpenStudy (nurali):

−2(x + 3) = −2(x + 1) − 4 -2x-6=-2x-2-4 -2x-6=-2x-6 -2x+2x=-6+6 0=0 All real numbers is the solution.

OpenStudy (calculusxy):

Isn't 1 supposed to be the answer? Since 1 is already in the other side of the equation.